OpenStudy (anonymous):

Find the equation of the line passing through the points A(3, −3) and B(−4, 11)? y = −6x + 3 y = −2x + 3 y = 9x −3 y = 4x − 9 i got the last one am i right?

OpenStudy (radar):

You have been provided with two points You can use the formula m=(y2-y)/(x2-x)

OpenStudy (anonymous):

option 2 is correct.use x-x1/x2-x1=y-y1/y2-y1

OpenStudy (radar):

m=-2

OpenStudy (anonymous):

you can directly find the equation of line :- as y-11=(slope)(x-(-4)) where slope is (11-(-3))/(-4-3) = -2 y-11 = -2(x+4) y-11= -2x -8 y + 2x = 3

OpenStudy (anonymous):

i.e y= -2x + 3

OpenStudy (radar):

The only option that has a slope it the 2nd one wher y=-2x+3

OpenStudy (anonymous):

thanks (: